if its 2:36 what time will it be in 45 mins
Answer by asian.go97(1)

36mins + 45mins = 81 minutes. One hour consists of 60 minutes so, 81 - 60 = 21. If the hour was 2, then add 1. It is 3:21
Another way to solve that would be subtract 36 mins from 60 (one hour): 60 - 36 = 24 minutes until 3:00
then subtract that from 45: 45 - 24 = 21 minutes.
Answer: 3:21